OpenStudy (anonymous):

What is the remainder when (3x^4 + 2x^3 – x^2 + 2x – 19) ÷ (x + 2) ? 0 10 5 15

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

remainder of 5?

OpenStudy (amoodarya):

(3x^4 + 2x^3 – x^2 + 2x – 19) =(x+2) g(x) +R if we put x=-2 then will have R easily

OpenStudy (amoodarya):



OpenStudy (amoodarya):

R=5

OpenStudy (anonymous):

sweet thanks man :)